# Algorithms for Dynamic Computational Geometry with Applications

### v1

@@ Abstract
Most of the literature of computational geometry concerns geometric properties of sets of static points. M.J. Atallah introduced “dynamic computational geometry,” concerned with both momentary and long-term geometric properties of sets of moving point-objects. This area of research seems to have been dormant recently. The current paper examines new problems in dynamic computational geometry.
Key words and phrases: dynamic computational geometry, piecewise defined continuous function, analysis of algorithms
@@ 1 Introduction
M.J. Atallah’s paper [1] presented efficient sequential algorithms for a variety of problems in computational geometry concerned with systems of point–objects in motion in Euclidean space ${\mathbb{R}}^{d}$; it was assumed for each object that its motion was described in each coordinate by known polynomials of time. Parallel versions of these algorithms were studied in [2, 3, 4]. In the current paper, we introduce efficient algorithms for additional problems in dynamic computational geometry, and discuss applications.
@@ 2.2 On more general motion
for appropriate constants $R_{1},R_{2},a,\theta_{0},x_{0},y_{0}$, where $t$ is the time variable. Note that all coordinates have the same period, and the $x$ and $y$ coordinates describe a planar ellipse (a circle if $R_{1}=R_{2}$), with constant angular velocity.
@@ 2.3 Piecewise defined functions
Among the problems studied in [1, 2, 3, 4] are
  • the Nearest Neighbor Problem: which pair of objects are nearest, as a function of time? and
  • the Collision Problem: when will some pair of objects collide?
Another important safety problem, which we call the “Too Close Problem”: when is the earliest, if ever, an object $q_{i}$ is too close to one or more other objects $q_{j}$?
@@
If for any index $j\neq 1$, inequality (2) holds for $i=1$ and $t=0$, then $t=0$ is the earliest time of a “too close” occurrence. Otherwise, at $t=0$, we have $d(P_{1},P_{j})>\min\{q_{1},q_{j}\}$ for all $j>1$, and since the trajectory functions $f_{i}$ are all continuous, any instances of inequality (2) occur earliest when $d(P_{1},P_{j})=\min\{q_{1},q_{j}\}$. Thus, Algorithm 3.1 considers these possibilities.
@@
Let $Q$ be a set of $n$ point-objects as in section 2.1. Assume for each $i$ we know $g_{i}$, the maximum safe distance between $q_{i}$ and any other $q_{j}$. We solve the Too Close Problem for $q_{1}$ as follows, where the return values of the algorithm are in the variables $minTime$ and $earliestIndices$.
@@
Thus, the algorithm has a running time of $\Theta(n)$. This is optimal, since each member of $Q$ must be considered. ∎
@@
This problem can be solved by an algorithm similar to Algorithm 3.1. The sequential execution time is therefore an optimal $\Theta(n)$.
@@
We are particularly interested in the first predictable occurrence of, say, $q_{1}$ and 2 other point-objects being (nearly) aligned, as perhaps we can call for at least one of them to change its trajectory before this occurs.
@@ 5.1 Preliminaries
In $\Theta(1)$ time, we can pick $t^{\prime}$ such that $u<t^{\prime}<v$ (e.g., $t^{\prime}$ could be the midpoint $(u+v)/2$ of $[u,v]$). By assumption, $f(t^{\prime})\neq 0$. Since there are no zeroes of $f(t)$ on $[u,v]$, the continuity of $f$ implies
@@
Then $r\leq v\leq s$. Since we regard $s$ as bounded, this takes $\Theta(1)$ time.
@@
Use Lemma 5.1 to determine the function $F$ of $pieceInterval_{a}$ in $\Theta(1)$ time.
Thus, $(pieceInterval_{a},pieceFunction_{a}):=([t_{a},t_{a+1}],F)$
$previousLeft:=a$   /* prepare for next interval */
leftleft :=a+1:=a+1    /* prepare for next interval */
@@
$pieceInterval_{previousLeft}:=pieceInterval_{previousLeft}\cup[t_{left},t_{right}]$
@@
If $t_{0}$ is not an endpoint of a piece of $F$ then the function of $F$ does not change its piece at $t_{0}$, i.e., for some $\varepsilon>0$, $F$ is identically equal to some $F_{i}$ on $t_{0}-\varepsilon,t_{0}+\varepsilon]$. Similarly, if $t_{0}$ is not an endpoint of a piece of $G$ then the function of $G$ does not change its piece at $t_{0}$, i.e., for some $\varepsilon>0$, $G$ is identically equal to some $G_{j}$ on $t_{0}-\varepsilon,t_{0}+\varepsilon]$. Therefore, FGF-G is identically equal to $F_{i}-G_{j}$ on $t_{0}-\varepsilon,t_{0}+\varepsilon]$. This establishes our assertion. ∎
@@ 5.2 Solving the 3 aligned problem
(Suggested by Lemma 2.5 of [2].) Let $F,G:[0,M]\to{\mathbb{R}}$ be continuous functions. Let $F$ be described by pieces based on the members of ${\cal F}:=\{f_{i}\}_{i=1}^{k}$. Let $G$ be described by pieces based on the members of ${\cal G}:=\{g_{j}\}_{j=1}^{m}$. Suppose for each pair $(h_{1},h_{2})$ of distinct members of ${\cal F}\cup{\cal G}$, the graphs of $h_{1}$ and $h_{2}$ have at most $s$ points of intersection in $[0,M]$. Then each of the functions $F+G$, FGF-G, and $F\cdot G$ has at most $k+m$ switchpoints, hence at most $k+m+1$ pieces based on, respectively,
@@
and these pieces can be computed in $\Theta(k+m)$ time. If $k,m$, and $s$ are regarded as bounded, the time estimate reduces to $\Theta(1)$.
@@
Let $\cal G^{\prime}$} be the union of switchpoints of members of $\cal G$} on $[0,M]$. be
@@
For each piece $\cal P$ of $F$ and each piece $\cal Q$ of $G$, use Lemma 5.9 to compute the pieces of FGF-G on the intersection of the interval of $\cal P$ and the interval of $\cal Q$ in $O(1)$ time. Since there are at most $(k+m)s+1$ switchpoints of FGF-G, the function $F_{G}$ has at most $(k+m)s+1$ pieces based on the members of $\{f_{i}-g_{j}\mid f_{i}\in{\cal F},g_{j}\in{\cal G}\}$, on $[0,M]$. This step executes in $O((k+m)s+1)$ time.
Clearly, this algorithm runs in $O((k+m)s+1)$ time.
@@
$\{\pm f_{i}\pm g_{j}\pm h_{u}\mid 1\leq i\leq k,~1\leq j\leq m,~1\leq u\leq r\}$
and all these switchpoints can be computed in $\Theta(a+b+c)$ time. If $k,m$,and $r$ are regarded as bounded, the time estimate reduces to $\Theta(1)$.
@@
It seems appropriate to consider a solution to the Too Close problem as yielding a solution to the 3-aligned problem.
@@
There are at most $s$ solutions of $p_{k}(f_{i}(t))-p_{k}(f_{j}(t))=0$. Let zeroes in [0,M] of $p_{k}(f_{i}(t))-p_{k}(f_{j}(t))$ be $Z_{k}=\{t_{k,1}.\ldots,t_{k,u_{k}}\}$, where $0\leq u_{k}\leq s$. Then the pieces of $D_{i,j}(t)$ have at most $ds$ switchpoints in $[0,M]$ at which some $k^{th}$ component switches its function. Therefore, $D_{i,j}(t)$ has at most $ds+1$ pieces based on $\{f_{i},f_{j}\}$.
@@
$q_{1}$ is approximately collinear with, and between, $q_{i}$ and $q_{j}$, if for sufficiently small $\varepsilon>0$,
$|f_{i}(t)-f_{j}(t)|+\varepsilon\geq|f_{i}(t)-f_{1}(t)|+|f_{1}(t)-f_{j}(t)|.$ (3)
@@
    $\varepsilon=\sum_{k=1}^{d}|p_{k}[f_{i}(t)-f_{1}(t)]|+\sum_{k=1}^{d}|p_{k}[f_{1}(t)-f_{J}(t)]|-\sum_{k=1}^{d}|p_{k}[f_{i}(t)-f_{j}(t)]|.$ (5)
    In this case, Corollary 5.7 and Corollary 5.11 yield that equation  (5) has at most  $3ds$ switchpoints and the pieces of the right side of (5), based on the members of
@@
By computing, for all index pairs $(i,j)$ such that $1<i<j\leq n$, $f_{1}(0)$, $f_{i}(0)$, and $f_{j}(0)$, we can determine from (4) any (if existing) solution occurs at $t=0$. Since there are $\Theta(n^{2})$ such pairs $(i,j)$, this takes $\Theta(n^{2})$ sequential time.
If no such solution exists, we set $t_{0}:=\infty$, and proceed as follows.
@@
Else /* min not found at 0 for current $i,j$ */
@@
End If no Too Close solution
Since each of $a(t),b(t),c(t)$ has at most $ds$ switchpoints, computation of their descriptions requires $\Theta(1)$ time. It follows that the worst-case sequential running time of the algorithm is $\Theta(n^{2})$, which is optimal since there are $\Theta(n^{2})$ pairs $(i,j)$ that must be considered in the worst case.

### v2

@@ Abstract
Most of the literature of computational geometry concerns geometric properties of sets of static points. M.J. Atallah introduced “dynamic computational geometry,” concerned with both momentary and long-term geometric properties of sets of moving point-objects. This area of research seems to have been dormant recently. The current paper examines new problems in dynamic computational geometry. Solutions are given for the sequential and coarse-grained multicomputer (CGM) models of computation.
Key words and phrases: dynamic computational geometry, piecewise defined continuous function, analysis of algorithms, coarse-grained multicomputer
@@ 1 Introduction
M.J. Atallah’s paper [1] presented efficient sequential algorithms for a variety of problems in computational geometry concerned with systems of point–objects in motion in Euclidean space ${\mathbb{R}}^{d}$; it was assumed for each object that its motion was described in each coordinate by known polynomials of time. Parallel versions of these algorithms were studied in [2, 3, 6, 7]. Another paper concerned with geometric properties of moving objects is [10]. The term “dynamic” is also applied to the study of a system of objects in which, over time, objects are inserted or deleted from the system, but this is not what is studied in the current paper.
In the current paper, we introduce efficient sequential and coarse grained parallel algorithms for additional problems in dynamic computational geometry (as introduced in Atallah’s paper [1]), and discuss applications.
@@ 2.2 On more general motion
for appropriate constants $R_{1},R_{2},a,\theta_{0},x_{0},y_{0}$, where $t$ is the time variable. Note that both coordinates have the same period, and the $x$ and $y$ coordinates describe a planar ellipse (a circle if $R_{1}=R_{2}$), with constant angular velocity.
@@ 2.3 Piecewise defined functions
2.4 Coarse grained multicomputer
Material in this section is largely quoted or paraphrased from [5].
The coarse grained multicomputer (CGM) was described in [8] as a model of parallel computation capable of processing a great deal of data with relatively few processors. A $CGM(n,p)$ has $p$ processors operating on $\Theta(n)$ data. Every processor has $\Omega(n/p)$ memory cells, each of $\Theta(\log n)$ bits. “Coarse grained” means the size $\Omega(n/p)$ of local memory is “considerable larger” than $\Theta(1)$; customarily, this is taken to mean $n/p\geq p$, i.e., $n\geq p^{2}$. We use this conventionin the current paper. Processors may share memory or may be arranged in some interconnection network.
Processors are indexed from 1 to $p$, and each processor “knows” its index.
We regard a $CGM(n,p)$ as a connected graph $G$ in which vertices are processors. If the processors are in an interconnection network then edges are communication links between directly connected processors. If the processors share memory then we regard $G$ as a complete graph, i.e., every pair of processors is regarded as joined by a communication link.
Processors were assumed in [8] to communicate data among themselves through sorting operations, but we will use the assumption of later papers that any pair of adjacent processors may exchange a unit of data in $\Theta(1)$ time. Similarly, in a shared memory system, we assume any pair of processors may exchange a unit of data in $\Theta(1)$ time.
We will use the following.
Theorem 2.5.
[4] A unit of data can be broadcast in a $CGM(n,p)$ from its source processor to all processors in $O(p)$ time.
A set of values $S$ distributed among the processors of a parallel computer $G$ is said to be gathered to a processor $P$ of $G$ when all the values of $S$ are copied to $P$.
Theorem 2.6.
[4] Let $S$ be a set of $N$ elementary data items distributed among the processors of a $CGM(n,p)$ $G$ such that $N=\Omega(p)$ and $N=O(n/p)$. $S$ can be gathered to any processor of $G$ in optimal $\Theta(N)$ time.
Theorem 2.7.
[4] Let $X=\{x_{i}\}_{i=1}^{n}$ be a set of elementary comparable data items distributed one per processor in a $CGM(n,p)$. The value of a semigroup operation (e.g.,  $\min\{x\in X\}$) can be computed and broadcast to all processors in optimal $\Theta(n/p)$ time.
@@ 2.4 Coarse grained multicomputer
Problems studied in [1, 2, 3, 6] that may be interpreted as concerned with safety:
  • The Nearest Neighbor Problem: which pair of objects are nearest, as a function of time?
  • The Collision Problem: when (if ever) will some pair of objects collide?
Another important safety problem, which we call the “Too Close Problem”: when is the earliest, if ever, the object $q_{1}$ is too close to one or more other objects $q_{j}$?
@@
If for any index $j\neq 1$, inequality (2) holds at $t=0$, then $t=0$ is the earliest time of a “too close” occurrence. Otherwise, at $t=0$, we have $d(P_{1},P_{j})>\min\{q_{1},q_{j}\}$ for all $j>1$, and since the trajectory functions $f_{i}$ are all continuous, any instances of inequality (2) occur earliest when $d(P_{1},P_{j})=\min\{q_{1},q_{j}\}$.
3.1 Sequential solution
@@
Let $Q$ be a set of $n$ point-objects as in section 2.1. Assume for each $i$ we know $f_{i}$, the function that describes the motion of $q_{i}$; and $g_{i}$, the minimum safe distance between $q_{i}$ and any other $q_{j}$. We solve the Too Close Problem for $q_{1}$ as follows, where the return values of the algorithm are in the variables $minTime$ and $earliestIndices$.
@@
Thus, the algorithm has a running time of $\Theta(n)$. This is worst-case optimal, since in the worst case, each member of $Q$ must be considered. ∎
Remark 3.3.
Note if all we want is the time of the first instance of too-closeness, and not the index set of the too-close objects, we can change the For loop to a loop governed by “While $minTime>0$ and $j\leq n$”. In this case we have a best-case running time of $\Theta(1)$, realized in the case $|S_{2}(0)|\leq d_{2}$.
3.2 CGM solution
We give a CGM solution that is only concerned with finding the earliest instance, if one exists, of two members of $Q$ being too close.
Algorithm 3.4.
Let $Q$ be a set of $n$ point-objects as in section 2.1. Assume each of the $p$ processors has descriptions of $\Theta(n/p)$ of the pairs $(f_{i},g_{i})$. Without loss of generality, processor $P_{j}$ has descriptions of the records in the set $\left\{(f_{i},g_{i})\}_{i=(j-1)n/p+1}^{jn/p}\right\}$. We solve the Too Close Problem for $q_{1}$ as follows, where the return values of the algorithm are in the variables $minTime$ and $earliestIndices$.
  1. 1.
    Broadcast the pair $(f_{1},g_{1})$ from processor $P_{1}$ to all processors.
  2. 2.
    Use Algorithm 3.1 so that each processor $P_{i}$ solves its portion of the problem, finding the earliest time $T_{i}$ that any of $\{q_{i}\}_{i=(j-1)n/p+1}^{jn/p}\}$ ($i\neq 1)$, if any, is too close to $q_{1}$.
  3. 3.
    Gather $\{T_{i}\}_{i=1}^{p}$ to $P_{1}$.
  4. 4.
    Processor $P_{1}$ computes $m=\min\{T_{i}\}_{i=1}^{p}$ in $\Theta(p)$ time and notes the set $Y$ of indices at which the minimum occurs. I.e., $j\in Y$ if and only if $T_{j}=m$.
  5. 5.
    If desired, broadcast $m$ from $P_{1}$ to all processors.
Theorem 3.5.
Algorithm 3.4 runs in optimal $\Theta(n/p)$ time.
Proof.
  1. 1.
    By Theorem 2.5, the broadcast requires $O(p)$ time.
  2. 2.
    By Theorem 3.1, the local partial solutions require parallel $\Theta(n/p)$ time.
  3. 3.
    By Theorem 2.6, the gather step requires $O(p)$ time.
  4. 4.
    The minimum computed in $P_{1}$ requires $\Theta(p)$ sequential time.
  5. 5.
    By Theorem 2.5, the broadcast requires $O(p)$ time.
Thus the algorithm runs in $\Theta(n/p)$ time.
The worst-case optimality of this running time follows from the worst-case optimality of $\Theta(n)$ as the running time of the sequential solution. ∎
Remark 3.6.
If we wish to modify Algorithm 3.4 to return the index set of the Too Close objects at the earliest instance of Too-closeness, notice that in the worst case, all members of $Q\setminus\{q_{1}\}$ could be too close to $q_{1}$ at that time, so the index set would have cardinality n1n-1, too large for the application of Theorem 2.6. We can obtain such an algorithm with the same asymptotic running time as Algorithm 3.4 by not gathering the index set.
@@ 3.2 CGM solution
This problem can be solved by algorithms similar to Algorithm 3.1 and 3.4. The sequential execution time is therefore a worst-case optimal $\Theta(n)$, and the execution time for a $CGM(n,p)$ is therefore a worst-case optimal $\Theta(n/p)$.
@@
We are particularly interested in the first predictable occurrence of, say, $q_{1}$ and 2 other point-objects being (nearly) aligned, as perhaps we can call for at least one of them to change its trajectory before this occurs. We give a worst-case optimal sequential solution for this problem.
@@ 5.1 Preliminaries
In $\Theta(1)$ time, we can pick $t^{\prime}$ such that $u<t^{\prime}<v$ (e.g., $t^{\prime}$ could be the midpoint $(u+v)/2$ of $[u,v]$). By assumption, $f(t^{\prime})\neq 0$. Since there are no zeroes of $f(t)$ on the interval $(u,v)$, the continuity of $f$ implies
@@
    Then $1\leq v\leq s$. Since we regard $s$ as bounded, this takes $\Theta(1)$ time.
@@
  • Use Lemma 5.1 to determine the function $F$ of $pieceInterval_{a}$ in $\Theta(1)$ time. Thus,
    $(pieceInterval_{a},pieceFunction_{a}):=([t_{a},t_{a+1}],F).$
  • $(previousLeft,Left):=(a,a+1)$   /* prepare for next interval */
@@
    $pieceInterval_{previousLeft}:=$
          $pieceInterval_{previousLeft}\cup[t_{left},t_{right}]$
@@
If $t_{0}$ is not an endpoint of a piece of $F$ then the function of $F$ does not change its piece at $t_{0}$, i.e., for some $\varepsilon_{1}>0$, $F$ is identically equal to some $F_{i}$ on $t_{0}-\varepsilon_{1},t_{0}+\varepsilon_{1}]$. Similarly, if $t_{0}$ is not an endpoint of a piece of $G$ then the function of $G$ does not change its piece at $t_{0}$, i.e., for some $\varepsilon_{2}>0$, $G$ is identically equal to some $G_{j}$ on $t_{0}-\varepsilon_{2},t_{0}+\varepsilon_{2}]$. Therefore, FGF-G is identically equal to $F_{i}-G_{j}$ on $t_{0}-\varepsilon_{3},t_{0}+\varepsilon_{3}]$, where $\varepsilon_{3}=\min\{\varepsilon_{1},\varepsilon_{2}\}$. This establishes our assertion. ∎
@@ 5.2 Solving the 3 aligned problem
(Suggested by Lemma 2.5 of [2].) Let $F,G:[0,M]\to{\mathbb{R}}$ be continuous functions. Let $F$ be described by pieces based on the members of ${\cal F}:=\{f_{i}\}_{i=1}^{k}$. Let $G$ be described by pieces based on the members of ${\cal G}:=\{g_{j}\}_{j=1}^{m}$. Suppose for each pair $(h_{1},h_{2})$ of distinct members of ${\cal F}\cup{\cal G}$, the graphs of $h_{1}$ and $h_{2}$ have at most $s$ points of intersection in $[0,M]$. Then each of the functions $F+G$, FGF-G, and $F\cdot G$ has at most $(k+m)s$ switchpoints, hence at most $(k+m)s+1$ pieces based on, respectively,
@@
and these pieces can be computed in $O(kms^{2}+(k+m)s)$ time. If $k,m$, and $s$ are regarded as bounded, the time estimate reduces to $\Theta(1)$.
@@
Let $\cal G^{\prime}$} be the union of switchpoints of members of $\cal G$} on $[0,M]$,
@@
For each pair $(\cal P,\cal Q)$ of pieces of $F$ and of $G$, respectively, use Lemma 5.9 to compute the pieces of FGF-G on the intersection of the interval of $\cal P$ and the interval of $\cal Q$ in $O(1)$ time. There are at most $kms^{2}$ such pairs and at most $(k+m)s+1$ switchpoints of FGF-G. The function FGF-G has at most $(k+m)s+1$ pieces based on the members of $\{f_{i}-g_{j}\mid f_{i}\in{\cal F},g_{j}\in{\cal G}\}$, on $[0,M]$. This step executes in $O((k+m)s+1)$ time.
Clearly, this algorithm runs in $O(kms^{2}+(k+m)s+1)$ time.
@@
$\{\pm f_{i}\pm g_{j}\pm h_{u}\mid 1\leq i\leq a,~1\leq j\leq b,~1\leq u\leq c\}.$
If $k,m,r,s$ are regarded as bounded, then all these switchpoints can be computed in $\Theta(1)$ time.
@@
There are at most $s$ solutions of $p_{k}(f_{i}(t))-p_{k}(f_{j}(t))=0$. Let zeroes in [0,M] of $p_{k}(f_{i}(t))-p_{k}(f_{j}(t))$ be $Z_{k}=\{t_{k,1}.\ldots,t_{k,u_{k}}\}$, where $0\leq u_{k}\leq s$. Then the pieces of $D_{i,j}(t)$ have at most $ds$ switchpoints in $[0,M]$ at which some $k^{th}$ component switches its base function. Therefore, $D_{i,j}(t)$ has at most $ds+1$ pieces based on $\{\pm[p_{k}(f_{i})-p_{k}(f_{j})]\}_{k=1}^{d}$.
@@
$q_{1}$ is approximately collinear with, and “between”, $q_{i}$ and $q_{j}$, if for sufficiently small $\varepsilon>0$,
$|f_{i}(t)-f_{j}(t)|+\varepsilon\geq|f_{i}(t)-f_{1}(t)|+|f_{1}(t)-f_{j}(t)|$ (3)
Similar inequalities apply for approximate collinearity with either $q_{i}$ “between” $q_{1}$ and $q_{j}$, or $q_{j}$ “between” $q_{1}$ and $q_{i}$.
@@
    $\varepsilon=\sum_{k=1}^{d}|p_{k}[f_{i}(t)-f_{1}(t)]|+\sum_{k=1}^{d}|p_{k}[f_{1}(t)-f_{J}(t)]|-\sum_{k=1}^{d}|p_{k}[f_{i}(t)-f_{j}(t)]|$ (5)
    or similarly for the other “between” possibilities. In this case, Corollary 5.7 and Corollary 5.11 yield that equation  (5) has at most  $3ds$ switchpoints and the pieces of the right side of (5), based on the members of
@@
In the following algorithm, we compute, for all index pairs $(i,j)$ such that $1<i<j\leq n$, $f_{1}(0)$, $f_{i}(0)$, and $f_{j}(0)$. This lets us determine from (4) any (if existing) solution occurs at $t=0$. If no such solution exists, we test for another solution.
@@
Else /* min not found at $t=0$ for current $i,j$ */
@@
Theorem 5.16.
Algorithm 5.15 runs in worst-case optimal $\Theta(n^{2})$ time and in best-case $\Theta(1)$ time..
Proof.
Since each of $a(t),b(t),c(t)$ has at most $ds$ switchpoints, each computation of candidatescandidates requires $\Theta(1)$ time. Since in the worst case there are $\Theta(n^{2})$ pairs $(i,j)$ that must be considered, it follows that the worst-case sequential running time of the algorithm is $\Theta(n^{2})$, which is optimal.
In the best case, we discover a solution at $t=0$ for $i=2$ and $j=3$, so that the While loop body is executed just once. Thus, in this case, $\Theta(1)$ time is used. ∎